how to calculate the formular as below :
1). 2X+3Y=21
2). X/2+Y/3=19/6
Pls help me to find out the best way to solve the problem, TQ

Answer by RAY100(1637)

remember three ways to solve linear equations, 1) graph, 2) substitution, 3)elimination
on this one let's try elimination
,
1) 2x+3y=21
2) x/2+y/3=19/6
,
to eliminate fractions let's multiply (2) by 6
,
2a) 3x +2y =19
,
Comparing 1) and 2a), we might multiply, 1) by 3,,and 2a) by 2,, to get common coefficients
'
1) 6x +9y =63
2a) 6x +4y =38
,
Now let's subtract 1)-2a)
,
5y=25
y=5,,,,,ANSWER
,
substitute into 1)
,
2x+3(5)=21
2x=21-15=6
x=3,,,,,,ANSWER
,
checking
1) 2(3)+3(5)=21
6+15=21,,,ok
2) (3)/2 +(5)3 = 19/6
1.5+1.667=3.167
3.167=3.167 ,,,ok